ILovePuppiesLol (ilovepuppieslol):

Fun Math Challenge Question! 2+3=8, 3+7=27, 4+5=32, 5+8=60, 6+7=72, 7+8=?? Rules are you can't use the internet ;)

OpenStudy (truwhovian11):

15

ILovePuppiesLol (ilovepuppieslol):

nope lmao @truwhovian11

ILovePuppiesLol (ilovepuppieslol):

its more complex than that @truwhovian11

OpenStudy (anonymous):

Hm...

OpenStudy (anonymous):

I'm seeing something with a pattern

OpenStudy (anonymous):

3, 5, and 7. Does it have naything to do with it?

OpenStudy (k12awesomeness):

I've done stuff like this before... it's really easy, the answer for 7+8 = 127

OpenStudy (k12awesomeness):

I think so anyway

OpenStudy (k12awesomeness):

is that the correct answer @ILovePuppiesLol

ILovePuppiesLol (ilovepuppieslol):

there are multiple methods to get the correct answer but it is not 127

OpenStudy (anonymous):

Oh! I think I got it. 98?

OpenStudy (k12awesomeness):

kk

ILovePuppiesLol (ilovepuppieslol):

yes it is 98, how did u get the answer?

OpenStudy (anonymous):

For each number you add. You multiply it by 1 less than that number. First I was trying to multiply by just 2, then just 3.. But that didn't work. Then I tried multiplying by different numbers. And eventually got that.

ILovePuppiesLol (ilovepuppieslol):

can u type it out?

ILovePuppiesLol (ilovepuppieslol):

how u got each solution?

ILovePuppiesLol (ilovepuppieslol):

nice job btw ;)

OpenStudy (anonymous):

First line: 2-1 = 1, 3-1 = 2. Then: 2*1 + 3 *2 = 8 Then say for 2nd to last line is 6*5 + 7*6 = 30 + 42 = 72.. Then last line is same thing, which is 98.